OpenStudy (anonymous):

There are two numbers with the following properties. 1) The second number is 3 more than the first number. 2) The product of the two numbers is 9 more than their sum. Which of the following represents possible values of these two numbers? A) -6,-3 B) -4,-1 C) -1,4 D) -3,6

OpenStudy (anonymous):

please help:)

OpenStudy (anonymous):

A or B

OpenStudy (anonymous):

Its B

OpenStudy (anonymous):

explain please

OpenStudy (anonymous):

Yep. B. -4 plus 3 is -1. And -4 times -1 is 4, which is nine more than (-4)+(-1) which is -5.

Directrix (directrix):

@bandicoot12 Reminder: OpenStudy values the Learning process - not the ‘Give you an answer’ process •Don’t post only answers - guide the asker to a solution.

OpenStudy (anonymous):

call the first number x and the second number y so first statement: The second number is 3 more than the first number. so it is y = 3 + x 2nd statement: The product of the two numbers is 9 more than their sum. so it is xy = 9 + x + y so just solve for one variable buy substiting in y x(3+x) = 9 + x + 3 + x 3x + x^2 = 12 + 2x x + x^2 - 12 = 0 factor this and get (x+4)(x-3) = 0 so x is either -4 or 3 substitute in y = -4 + 3 = -1 and y = 3 + 3 = 6 so your possible solutions are (-4, -1) or (3,6) so its B